OpenStudy (anonymous):

5 over square root of 2 .

OpenStudy (anonymous):

the answer is squaroot pf 5!

OpenStudy (anonymous):

\[5/\sqrt{2}\] needs to be rationalized because we cannot leave a radical in the denominator of a fraction. \[(5/\sqrt{2} )* (\sqrt{2}/\sqrt{2})\] So we multiply by sqrt2 over sqrt2 which equals 1. It allows us to change how it looks without changing the value. and we get \[5\sqrt{2}/2\]